Find a quadratic function of the form y=ax^2 that passes through the point ​(​3,-27).
Answer by math_tutor2020(3816)


y = a*x^2
-27 = a*3^2 .... plug in (x,y) = (3,-27)
-27 = a*9
-27 = 9a
9a = -27
9a/9 = -27/9 ..... divide both sides by 9
a = -3

Answer: y = -3x^2